OpenStudy (anonymous):

Benny has a total of 64 quarters and pennies in his pocket. These coins total $4.00. How many pennies are in Benny's pocket?

OpenStudy (anonymous):

Let me help you

OpenStudy (anonymous):

Let the no of quarters be x and no of pennies be y

OpenStudy (anonymous):

what are the toal no of coins in his pocket?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

x+y=64=total coins

OpenStudy (anonymous):

total money in his pocket?

OpenStudy (anonymous):

4$

OpenStudy (anonymous):

if he has x quarters and y pennies.. the total money is 0.25x + 0.01y = 4

OpenStudy (anonymous):

You have two equations

OpenStudy (anonymous):

Dude onle a nerd would know this kinda stuff! I'm not a nerd, i just want to pass this stupid algebra test, I don't want to work anything out.

OpenStudy (anonymous):

Well its beyond me to try to exlain

OpenStudy (anonymous):

And you're just confusing me more, I already told you I don't understand any of this,

OpenStudy (anonymous):

I don't want an explaination, sorry but i don't. I get more than enough lectures at home, i dont need it from someone on a computer too, i just want an answer...

OpenStudy (anonymous):

x=14, y=50

OpenStudy (anonymous):

x+y=64